OpenStudy (anonymous):

Find the slope and the y-intercept of the equation y − 3(x − 1) = 0.

OpenStudy (anonymous):

slope = -3 and y-intercept = -3 slope = 3 and y-intercept = 3 slope = -3 and y-intercept = 3 slope = 3 and y-intercept = -3

OpenStudy (anonymous):

@iGreen @texaschic101

OpenStudy (ikram002p):

stander formula of line equation with slope / point is y= m x + b s.t m os the slope and b is y-intercept

OpenStudy (anonymous):

I know that

OpenStudy (ikram002p):

u have y − 3(x − 1) = 0. so y=3(x-1) y=3x +(-3 ) copare to this y= m x + b

OpenStudy (igreen):

Going with what @ikram002p said. You have y = 3x - 3 It's in the form of y = mx + b Where m = slope, and where b = y-intercept. Now do you know what they slope and y-intercept is? @usitdown

OpenStudy (anonymous):

No thats the question im asking

OpenStudy (anonymous):

The very top are the choices, but i dont know how to find the slope and y intercept

OpenStudy (igreen):

y = 3x - 3 ^ = Slope y = 3x - 3 ^ = y-intercept

OpenStudy (igreen):

So what's the slope and y-intercept? @usitdown

OpenStudy (anonymous):

Ohhhh Thanks

OpenStudy (igreen):

So what answer did you get? Which option?

OpenStudy (anonymous):

D

OpenStudy (igreen):

That's correct :)

OpenStudy (anonymous):

You are a big help @iGreen. If i knew you i would buy you some fried chicken

OpenStudy (igreen):

Lel